# The Scale for the Expansion of the Universe: From Local Structures to Cosmology

### v1

@@ Abstract
This study tackles the impact dark energy in different systems by a simple unifying formalism. We introduce a parameter space to compare gravity tests across all cosmic scales, using the McVittie spacetime (MCV), that connect spherically symmetric solutions with cosmological solutions. By analyzing invariant scalars, the Ricci, Weyl, and Kretschmann scalars, we develop a phase-space description that predicts the dominance of the Cosmological Constant. We explore three cases: (1) the local Hubble flow around galaxy groups and clusters, (2) spherical density distributions and (3) binary motion. Our results show that galaxy groups and clusters exhibit Kretschmann scalar values consistent with the Cosmological Constant curvature, indicating where dark energy dominates.
@@ I Introduction
Dark Energy (DE) is one of the most enigmatic puzzles in contemporary cosmology. Introduced to account for the detected accelerated expansion of the cosmos, DE is hypothesized to be an unidentified energy type that pervades all of space, generating repulsive effects to offset gravitational pull on large cosmic scales Peebles and Ratra (2003). Its identification, prompted by observations of Type Ia supernovae, revolutionized our comprehension of the universe’s structure and evolutionary trajectory Perlmutter et al. (1999). The concept of DE as the principal driver of cosmic acceleration has gained reinforcement from diverse observational data, such as measurements of the cosmic microwave background (CMB) and analyses of large-scale cosmic structures. Although DE is primarily examined in cosmological contexts, its impact also reaches smaller astrophysical environments, including galaxy clusters, particularly near transitional zones like the zero-velocity surface, where universal expansion shifts to localized gravitational collapse Pavlidou et al. (2014). This interplay between local and global DE effects provides a unique opportunity to probe its properties across a wide range of scales.
This study examines DE phenomena across cosmological and local scales using the MCV spacetime as a theoretical model. The MCV metric characterizes a spherically symmetric mass embedded in an expanding universe, serving as a versatile tool to analyze DE’s interplay between localized and cosmic regimes. By studying the spacetime’s curvature features, we pinpoint areas where DE’s effects dominate and assess their relevance for testing gravitational theories across different scales—from strong-field environments near dense objects to large-scale cosmological evolution, extending the work by Baker et al. (2015). This methodology bridges DE’s influence on cosmic dynamics, connecting its behavior in compact astrophysical settings to its role in shaping the universe’s large-scale structure.
This paper is organized as follows: Section II discuss the spacetime and its limiting cases. Section III discusses the geodesic equation for the MCV spacetime. Section IV discusses the expansion domination in different regimes. Section V discusses the results.
@@ II General Spacetime
A general framework for describing bound systems embedded in an expanding cosmological background is provided by the MCV metric McVittie (1933); Nandra et al. (2012). For a flat cosmological background, this metric takes the form:
$ds^{2}=-\left(1-\Phi-\frac{r^{2}H^{2}}{c^{2}}\right)c^{2}dt^{2}$
$-\frac{2rH}{\sqrt{1-\Phi}}dt,c,dr+\frac{dr^{2}}{1-\Phi}+r^{2}d\Omega^{2},$ (1)
where $\Phi\equiv 2GM/rc^{2}$ represents the gravitational potential, $H\equiv\dot{a}/a$ is the Hubble parameter and $\Omega$ is the angular part of the metric. Here, $M$ is the total mass of the structure, $G$ is the Newtonian gravitational constant, and $c$ is the speed of light. The metric reduces to the Friedmann-Lemaître-Robertson-Walker (FLRW) metric in the absence of a local mass ($\Phi=0$ and $M=0$), and to the Schwarzschild metric when the Hubble parameter vanishes ($H=0$), corresponding to a static, isolated mass. In the special case where $H=\text{Const}$ and $\Phi=0$, the metric simplifies to the de Sitter metric, describing a universe dominated by a cosmological constant. In Eq. (1), $r$ represents the physical spatial coordinate, which is related to the comoving coordinate by $\chi=r/a(t)$. By setting $M=0$ and using the comoving coordinate, we recover the standard flat FLRW metric, which describes a homogeneous and isotropic universe:
@@
Similarly, setting $H=0$ reduces the metric to the Schwarzschild metric. See Antoniou and Perivolaropoulos (2016) for a further discussion.
@@
To analyze the curvature properties of the MCV spacetime, we compute three key curvature scalars: the Ricci scalar $\mathcal{R}$, the Weyl scalar $\mathcal{C}$, and the Kretschmann scalar $\mathcal{K}$. These scalars provide a comprehensive description of the spacetime’s geometry, capturing the interplay between local gravitational effects and cosmic expansion. The Ricci scalar $\mathcal{R}$ reflects the overall curvature influenced by both the local mass and the cosmological expansion. The Weyl scalar $\mathcal{C}$ characterizes the tidal forces generated by the local mass, which dominate near compact objects. The Kretschmann scalar $\mathcal{K}$ combines contributions from both local and global curvature, offering a complete measure of the spacetime’s curvature invariants:
$\mathcal{R}=\frac{12H^{2}}{c^{2}}\left(1-\frac{3}{4}\frac{1+w}{\sqrt{1-\Phi}}\right),\quad\mathcal{C}=12\left(\frac{\Phi}{r^{2}}\right)^{2},$
$\mathcal{K}=\frac{12\Phi^{2}}{r^{4}}+\frac{24H^{4}}{c^{4}}\left(1-\frac{3}{2}\frac{1+w}{\sqrt{1-\Phi}}+\frac{9}{8}\frac{\left(1+w\right)^{2}}{1-\Phi}\right),$ (3)
@@
where $w\equiv-1-2\dot{H}/3H^{2}$ is the equation of state of the background universe. For a dark energy-dominated universe, $w=-1$; for a matter-dominated universe, $w=0$; and for a radiation-dominated era, $w=1/3$. To quantify the relative strength of local gravitational effects compared to cosmic expansion, we introduce the following dimensionless parameters:
$D_{H}\equiv\frac{c}{H},\quad r_{s}=\frac{2GM}{c^{2}}=r\Phi,$
$\kappa\equiv\frac{\Phi}{r^{2}}/\frac{2H^{2}}{c^{2}}=\frac{GM}{H^{2}r^{3}},\quad\tau\equiv\frac{1+w}{\sqrt{1-\Phi}}.$ (4)
Using these parameters, the curvature scalars can be expressed as:
$\mathcal{R}=\frac{12}{D_{H}^{2}}\left(1-\frac{3}{4}\tau\right),\quad\mathcal{C}=12\frac{r_{s}^{2}}{r^{6}},$
$\mathcal{K}=48\frac{r_{s}^{2}}{r^{6}}+\frac{24}{D_{H}^{4}}\left(1-\frac{3}{2}\tau+\frac{9}{8}\tau^{2}\right).$ (5)
@@
The MCV metric enables the investigation of how bound systems, such as galaxies or galaxy clusters, evolve within an expanding cosmological background. The dimensionless parameter $\kappa$ determines the ratio between these two parts and as we will see from different systems has different interpretation.
III Test Particle Geodesic
To describe binary motion in the presence of dark energy, we consider the low-energy limit of the geodesic equation over the metric (1), which gives:
$\ddot{r}=-\frac{GM}{r^{2}}+\frac{\ddot{a}}{a}r=-\frac{GM}{r^{2}}-\frac{(1+3w)}{2}H^{2}r.$ (6)
The first term represents the Newtonian gravitational attraction, while the second term accounts for the influence of dark energy, which acts as a repulsive force proportional to the Hubble parameter.

### v2

# Multiscale Cosmic Curvature: from Local Structures to Cosmology

@@ Abstract
This study tackles the impact Dark Energy (DE) in different systems by a simple unifying formalism. We introduce a parameter space that compare gravity across all cosmic scales, using the McVittie spacetime (McV) and connects spherically symmetric solutions with cosmological solutions. By analyzing the invariant scalars: the Ricci, Weyl, and Kretschmann scalars, we develop a phase-space description that predicts the dominance of the Cosmological Constant. We explore three cases: (1) the local Hubble flow around galaxy groups and clusters, (2) spherical density distributions and (3) binary motion. Our results show that the Kretschmann scalar of galaxy groups and clusters in their turnaround is $2\Lambda^{2}$ which is three times the Kretschmann scalar of the Cosmological Consonant. This indicates the DE domination in local structures.
@@ I Introduction
Dark Energy (DE) remains one of the most elusive components of modern cosmology. First introduced to explain the observed accelerated expansion of the universe, DE is thought to be a pervasive form of energy with negative pressure, capable of driving repulsive gravitational effects that counterbalance the attraction of matter on cosmic scales Peebles and Ratra (2003). The discovery of the unexpected dimming of distant Type Ia supernovae provided the first compelling evidence for this component, fundamentally reshaping our understanding of cosmic evolution and the large-scale structure of the Universe Perlmutter et al. (1999). Subsequent measurements of the cosmic microwave background (CMB)Aghanim et al. (2020a), baryon acoustic oscillations (BAO), and large-scale structure surveys have since reinforced DE’s role as a dominant constituent of the cosmic energy budgetDi Valentino et al. (2025).
While DE is commonly treated as a smooth background field, its influence on local dynamics—particularly in environments such as galaxy clusters and near zero-velocity surfaces—offers a complementary means of probing its properties Pavlidou and Fields (2005); Nasonova et al. (2011); Pavlidou and Tomaras (2014); Pavlidou et al. (2014); Tanoglidis et al. (2015, 2016); Lopes et al. (2018); Silbergleit and Chernin (2019); Bhattacharya and Tomaras (2021); Kim et al. (2020); Pavlidou et al. (2020); Paraskevas and Perivolaropoulos (2024); Del Popolo et al. (2020); Del Popolo and Chan (2020). Transitional region where cosmic expansion balances local gravitational attraction—are especially valuable for this purpose, serving as natural laboratories for testing how DE interacts with matter on intermediate scales Sandage (1986); Peirani and de Freitas Pacheco (2006); Peirani and Pacheco (2008); Peñarrubia et al. (2014); Sorce et al. (2016); Kim et al. (2020); Del Popolo et al. (2021); Del Popolo and Chan (2022); Aluri et al. (2023); Wagner and Benisty (2025).
A natural theoretical framework for exploring these multiscale effects is the McVittie (McV) spacetime McVittie (1933), which describes a spherically symmetric mass embedded in an expanding cosmological background Faraoni and Jacques (2007); Nolan (1998). The McV metric interpolates between Schwarzschild-like strong-field gravity Antoniou and Perivolaropoulos (2016) and FLRW cosmology, making it well suited for studying the interplay between local gravitational fields and cosmic expansion Nandra et al. (2012a, b). Curvature invariants such as the Ricci scalar, Kretschmann scalar, and Weyl invariants provide powerful tools for diagnosing this interplay Iihoshi et al. (2007); Baker et al. (2015); Moulin (2025). Their behavior can reveal regions where DE dominates, quantify tidal and shear effects, and trace how local overdensities modulate the effective strength of cosmic acceleration, extending analyses such as Baker et al. (2015).
In inhomogeneous regions, deviations from ideal FLRW behavior are encoded in curvature invariants, offering a diagnostic of DE-driven dynamics. For example, the Kretschmann scalar highlights the transition from mass-dominated to DE-dominated regimes, while Weyl curvature encodes the anisotropic imprint of local structures Baker et al. (2015). Linking these curvature measures to local kinematics bridges theoretical predictions with observations, from the dynamics of binaries Benisty et al. (2023, 2024) and dwarf galaxies in groups and clusters Karachentsev et al. (2009, 2007, 2002, 2013); McConnachie (2012); McConnachie and Venn (2020a, b); Tully et al. (2023) to cluster virialization and the universe’s large-scale expansion Gunn and Gott (1972); Peebles (1980); Pace et al. (2017). Comparisons between predicted DE-dominant regions and observed peculiar velocity fields allow us to test whether local expansion is consistent with a cosmological constant or indicates more complex DE behavior. With the advent of high-precision galaxy kinematic surveys and cosmological simulations Nelson et al. (2019a), it is now possible to map spatial variations in the local expansion rate and use them as a probe of DE–gravity interplay.
In this work, we study DE phenomena across cosmological and local scales within the McVittie framework. By computing curvature invariants and analyzing geodesic motion, we identify regimes where DE effects prevail over local gravity and assess their observational consequences for local Hubble flows. Our approach connects theoretical diagnostics with empirical data, providing a unified framework to test DE’s influence from compact astrophysical systems to the largest cosmological structures.
The paper is structured as follows: Section II introduces the McVittie spacetime and its limiting cases. Section III presents the geodesic equations in this setting. Section IV analyzes the regimes of expansion dominance using different datasets. Section V concludes with a discussion of results and their implications for future DE studies.
@@ II General Spacetime
II.1 The Curvature
A general framework for describing bound systems embedded in an expanding cosmological background is provided by the McV metric McVittie (1933); Nandra et al. (2012a). For a flat cosmological background, this metric takes the form:
$ds^{2}=-\left(1-\Phi-\frac{r^{2}H^{2}}{c^{2}}\right)c^{2}dt^{2}-\frac{2rH}{\sqrt{1-\Phi}}c\,dt\,dr$ (1)
$+\frac{dr^{2}}{1-\Phi}+r^{2}d\Omega^{2},$
where:
$\Phi\equiv\frac{2GM}{rc^{2}},\quad H\equiv\frac{\dot{a}}{a},$ (2)
represents the gravitational potential $\Phi$ and the Hubble parameter $H$ and $\Omega$ is the angular part of the metric. $M$ is the total mass of the structure, $G$ is the Newtonian gravitational constant and $c$ is the speed of light and $a$ is the scale factor of the spacetime. In Eq. (II.1), $r$ represents the physical spatial coordinate, which is related to the comoving coordinate by $\chi=r/a(t)$.
This choice of physical coordinates (as opposed to isotropic coordinates) is motivated by its direct physical interpretation for local systems and its straightforward relation to the comoving radius $\chi$, which naturally describes the expanding background in standard FLRW cosmology. The metric reduces to the FLRW metric in the absence of a local mass ($\Phi=0$ and $M=0$), and to the Schwarzschild metric when the Hubble parameter vanishes ($H=0$), corresponding to a static, isolated mass. In the special case where $H=\text{Const}$ and $\Phi=0$, the metric simplifies to the de Sitter metric, describing a universe dominated by a cosmological constant.
By setting $M=0$ and using the comoving coordinate, we recover the standard flat FLRW metric, which describes a homogeneous and isotropic universe:
@@ II.1 The Curvature
Similarly, setting $H=0$ reduces the metric to the Schwarzschild metric Antoniou and Perivolaropoulos (2016).
To investigate how local and cosmological effects combine in this spacetime, it is useful to compute the connection coefficients. For simplicity, we define the parameters:
$S\equiv\sqrt{1-\Phi},\quad D_{H}\equiv\frac{c}{H}\quad w\equiv-1-\frac{2\dot{H}}{3H^{2}}.$ (4)
where $D_{H}$ is the Hubble distance and $w$ is the equation of state of the Universe. For a DE-dominated universe, w=1w=-1; for a matter-dominated universe, $w=0$; and for a radiation-dominated era, $w=1/3$. Using these abbreviations, the Christoffel symbols can be written in an elegant way:
$\Gamma^{t}{}_{tt}$ $=\frac{2r^{2}-D_{H}^{2}\Phi}{2D_{H}^{3}S},\quad\Gamma^{t}{}_{tr}=\frac{\Phi-\tfrac{2r^{2}}{D_{H}^{2}}}{2rS^{2}},$
$\Gamma^{t}{}_{rr}$ $=\frac{1}{D_{H}S^{3}},\quad\Gamma^{t}{}_{\theta\theta}=\frac{r^{2}}{D_{H}S},\quad\Gamma^{t}{}_{\phi\phi}=\frac{r^{2}\sin^{2}\theta}{D_{H}S},$
$\Gamma^{r}{}_{tt}$ $=\frac{r^{3}}{D_{H}^{4}}+\frac{r}{2D_{H}^{2}}(1-S)-\frac{S^{2}\Phi}{2r}+\frac{3rS}{2D_{H}^{2}}(1+w),$
$\Gamma^{r}{}_{tr}$ $=\frac{D_{H}^{2}\Phi-2r^{2}}{2D_{H}^{3}S},\quad\Gamma^{r}{}_{rr}=\frac{D_{H}^{2}\Phi-2r^{2}}{2D_{H}^{2}rS^{2}},$
$\Gamma^{r}{}_{\theta\theta}$ $=r\left(\frac{r^{2}}{D_{H}^{2}}-S\right),\quad\Gamma^{r}{}_{\phi\phi}=\frac{r^{2}-D_{H}^{2}S^{2}}{D_{H}^{2}}r\sin^{2}\theta,$
$\Gamma^{\theta}{}_{r\theta}$ $=\Gamma^{\phi}{}_{r\phi}=\frac{1}{r},\quad\Gamma^{\theta}{}_{\phi\phi}=-\sin\theta\cos\theta,\quad\Gamma^{\phi}{}_{\theta\phi}=\cot\theta.$
Together with the radial and angular components, which reproduce the usual Schwarzschild and FLRW limits in the appropriate regimes. From these one constructs the Riemann tensor, whose independent nonzero components are:
$R_{\hat{t}\hat{r}\hat{t}\hat{r}}$ $=\frac{1}{D_{H}^{2}}-\frac{\Phi^{2}}{2r^{2}}\;+\;\frac{3}{2}\,(1+w)\,\frac{S}{D_{H}^{2}},$
$R_{\hat{t}\hat{\theta}\hat{t}\hat{\theta}}$ $=R_{\hat{r}\hat{\theta}\hat{r}\hat{\theta}}=\frac{1}{2}\!\left(\Phi-\frac{2r^{2}}{D_{H}^{2}}\right),$
$R_{\hat{r}\hat{\theta}\hat{\theta}\hat{t}}$ $=\frac{3}{2}\,(1+w)\,\frac{r\,c}{D_{H}^{3}}\;,$
$R_{\hat{r}\hat{\phi}\hat{\phi}\hat{t}}$ $=\;\frac{3}{2}\,(1+w)\,\frac{r\,c}{D_{H}^{3}}\;\,\sin^{2}\theta,$
$R_{\hat{\theta}\hat{\phi}\hat{\theta}\hat{\phi}}$ $=-\Big{(}\Phi+\frac{r^{2}}{D_{H}^{2}}\Big{)}\,\sin^{2}\theta.$
The above curvature components clearly interpolate between the Schwarzschild tidal terms and the cosmological background curvature. To understand the matter content that sources this geometry, we next evaluate the Einstein tensor $G_{\mu\nu}$. Its components take the form:
$G^{0}{}_{0}$ $=\frac{3}{D_{H}^{2}},\quad G^{1}{}_{0}=\frac{3rc}{D_{H}^{2}}(1+w),$
$G^{j}{}_{j}$ $=-\frac{3}{D_{H}^{2}}\left(1-\frac{1+w}{S}\right),$ (5)
with $j=1,2,3$ labeling the spatial components. The mixed component encodes a momentum flux proportional to $(1+w)$, vanishing in the cosmological constant case. The spatial components reveal anisotropic stresses induced by the local gravitational potential $\Phi$, signaling a departure from a pure perfect-fluid form in the presence of a compact object. Physically, the energy–momentum tensor thus consists of a homogeneous cosmological fluid driving the expansion, together with an inhomogeneous contribution due to the local mass. The combined source reproduces the hybrid McV geometry, which smoothly interpolates between the FLRW and Schwarzschild limits.
A distinctive feature of this effective matter source is that, unlike the strictly homogeneous FLRW case, the pressure acquires a nontrivial radial dependence through the potential $\Phi$. Close to the surface $r=2M/a(t)$ this pressure diverges, leading to violations of the strong energy condition Nolan (1998); Faraoni and Jacques (2007). These pathologies underline that the McV spacetime, while exact, is not supported by a physically realistic perfect fluid near the would-be black hole horizon.
The horizon structure itself is also subtle. In dust-dominated backgrounds ($w=0$) the McV horizon evolves dynamically, can fail to form smoothly, and may even bifurcate into multiple branches Nolan (1998); Faraoni and Jacques (2007). This peculiar behavior complicates the interpretation of how local Schwarzschild-like physics transitions into the cosmological FLRW regime. It should be kept in mind that extending the discussion into the near-horizon region would require caution, since violation of the energy condition and the dynamical horizon structure could substantially affect physical conclusions.
@@
To analyze the curvature properties of the McV spacetime, we compute three key curvature scalars: the Ricci scalar $\mathcal{R}$, the Weyl scalar $\mathcal{C}$, and the Kretschmann scalar $\mathcal{K}$. The Ricci scalar $\mathcal{R}$ reflects the overall curvature influenced by both the local mass and the cosmological expansion. The Weyl scalar $\mathcal{C}$ characterizes the tidal forces generated by the local mass, which dominate near compact objects. This scalar depends solely on the local mass because the FLRW background is conformally flat and thus has vanishing Weyl curvature, supported by Iihoshi et al. (2007) in Eq. (4.28) and Moulin (2025). The Kretschmann scalar $\mathcal{K}$ combines contributions from both local and global curvature, offering a complete measure of spacetime curvature invariants:
$\mathcal{R}$ $=\frac{12}{D_{H}^{2}}\left(1-\frac{3(1+w)}{4S}\right),\quad\mathcal{C}=12\left(\frac{\Phi}{r^{2}}\right)^{2},$
$\mathcal{K}$ $={12}\left(\frac{\Phi}{r^{2}}\right)^{2}+\frac{3}{D_{H}^{4}}\left(8-\frac{12(1+w)}{S}+\frac{9(1+w)^{2}}{S^{2}}\right),$ (6)
@@
To quantify the relative strength of local gravitational effects compared to cosmic expansion, we introduce the following physically motivated dimensionless parameters:
$\kappa\equiv\frac{\Phi}{r^{2}}\bigg{/}\frac{2H^{2}}{c^{2}}=\frac{GM}{H^{2}r^{3}},\quad\tau\equiv\frac{1+w}{\sqrt{1-\Phi}}.$ (7)
@@
$\kappa$ represents the ratio of local density ($\sim M/r^{3}$) to cosmological critical density ($3H^{2}/8\pi G$), while $\tau$ encodes the equation-of-state modification by gravitational potential. The McV metric enables the investigation of how bound systems, such as galaxies or galaxy clusters, evolve within an expanding cosmological background. The dimensionless parameter $\kappa$ determines the ratio between these two parts and as we will see from different systems has different interpretation.
III The Geodesic Equation
III.1 The Equations of Motion
The dynamics of a test particle moving solely under gravity around a central mass in a static spacetime can be described using the geodesic formalism. Restricting motion to the equatorial plane $\theta=\pi/2$, the Lagrangian corresponding to the metric (II.1) is
${\cal{L}}=\left[1-\Phi-\frac{r^{2}H^{2}}{c^{2}}\right]\dot{t}^{2}+\frac{2rH}{\sqrt{1-\Phi}}\dot{r}\dot{t}-\frac{\dot{r}^{2}}{1-\Phi}-r^{2}\dot{\phi}^{2},$ (8)
where overdots denote derivatives with respect to the proper time $\tau$. The Euler–Lagrange equations for $x^{\mu}=\{t,r,\phi\}$ yield the corresponding geodesics. In the case of stationary, spherically symmetric spacetimes such as the Schwarzschild metric first integrals of the $t$ and $\phi$ equations can be found directly, while the $r$ equation is often replaced by the normalization condition: $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=1$, (for timelike curves). This allows one to construct an energy-like equation for $r(\tau)$, from which $\ddot{r}$ may be obtained by differentiation. In contrast, the metric under consideration here depends on both $r$ and $t$, and thus the usual procedure cannot be directly applied. The radial geodesic equation retains an explicit $\ddot{r}$ term, which is essential for understanding the effective forces acting on the particle.
The geodesic equation for the azimuthal angle reduces to $l=\dot{\phi}r^{2}$, where $l$ is the particle’s specific angular momentum. Substituting this into the radial equation and rearranging gives  Nandra et al. (2012a, b):
$\ddot{r}=$ $\frac{l^{2}}{r^{3}}\left(1-\frac{3}{2}\Phi\right)-\frac{\Phi}{2}+rH^{2}-\frac{2r^{2}\dot{H}}{1-\Phi-r^{2}H^{2}}\dot{r}\dot{t}$
$+\frac{\dot{H}rS}{1-\Phi-r^{2}H^{2}}\left(1+\frac{l^{2}}{r^{2}}\right)+\frac{r\dot{H}\dot{r}^{2}}{S\left(1-\Phi-r^{2}H^{2}/c^{2}\right)}.$ (9)
Equation (9) provides the exact form of $\ddot{r}$, valid in the full relativistic regime. In the weak-field limit assuming:
$r_{s}\equiv\frac{GM}{c^{2}}\ll r\ll D_{H},\quad\Phi\ll 1,$ (10)
with $\dot{t}\approx 1$, $\dot{r}\approx 0$, we find the equation of motion to give:
$\ddot{r}=\frac{l^{2}}{2r^{3}}-\frac{GM}{r^{2}}-\frac{(1+3w)}{2}H^{2}r.$ (11)
The term $(1+3w)H^{2}/2$ is often written as $\equiv\ddot{a}/a$. An analytical solution for this equation of motion is studied in Baushev (2019) as integral terms. Here we connect the spherical collapse model with modifications close to the velocity surface as a solution.
Figure 1: The normalized velocity-distance relation from the analytical solution (17). The solution is scaled with $r_{0}$ and $r_{0}/t_{U}$, where $t_{U}$ is the age of the Universe. The turnaround $r_{0}$ is related to the enclosed mass. The colored lines corresponds to $e=1$. For different $e$ the area between the colored lines is also covered. The dashed line marks the asymptotic behavior for the unbound solution (which in the case of $H\neq 0$ corresponds to the Hubble flow).
III.2 Hubble Flow Around Virilized Structures
Figure 2: Left: The velocity vs. distance for an isolated pair from the IllustrisTNG simulation. Right: The radial velocity of galaxies as a function of their distance from the halo CoM. A green dashed line illustrates the theoretical Hubble Law with a slope of $70\,\text{km/s/Mpc}$, representing the expected outward velocity for galaxies due to the universe’s expansion.